# Connecting GRBs from Binary Neutron Star Mergers to Nuclear Properties of Neutron Stars

### v1

@@ III.2 Long and Short GRBs from BNS mergers
Our analysis assumes that the VLRNS population does not contribute to a sizeable fraction of cbGRBs for the reasons discussed earlier. However, as noted by Margalit et al. [61], some initially long-lived magnetars may evolve similarly to LRNSs, potentially powering an sbGRB. Therefore, we assess the implications of this scenario on our results by performing two additional simulations that consider the two extreme cases of the VLRNS population contributing fully to either the short or the long cbGRB components, depending on whether the VLRNSs collapse on a timescale of $\lesssim 1\,{\rm s}$ or $\sim 10\,{\rm s}$, respectively. We find that if VLRNSs contribute exclusively to sbGRBs, the transition with $a=1.2$ is still inconsistent with the data for $M_{\rm TOV}<2.5M_{\odot}$, while the $a=1.25$ threshold is only allowed for $M_{\rm TOV}\gtrsim 2.4M_{\odot}$. The higher value $a=1.3$ has the best consistency in the 2.2-2.4 $M_{\odot}$ TOV range of masses, while higher values of $a$ would require softer EoSs closer to 2.2 $M_{\odot}$. Conversely, if VLRNS contribute to the long cbGRB population only, we find that, within the 2.2-2.4 $M_{\odot}$ range of TOV masses, $a\lesssim 1.3$ is borderline consistent with the data, while both $a=1.35$ and $a=1.4$ have a broad range of consistency. These cases discussed above would constitute the most extreme departures from our fiducial model since either a smaller contribution from the VLRNS population or a mixed contribution to both the long and short cbGRBs would make the final estimates closer to those shown in Fig. 4. Therefore, our results of a relatively high transition mass $M_{\rm ls}\gtrsim 1.3M_{\rm TOV}$ are quite robust within the limits and assumptions of the unified GRB model and the available observational data to date.
@@ III.3 Comparison with numerical relativity
Our results are summarized in Fig. 5. The data shows clear evidence for a transition around ${\sim}1.3M_{\odot}$, consistent with the population synthesis results. The transition region is broad, spanning $1.3\lesssim M/M_{\rm TOV}\lesssim 1.4$, suggesting that there is additional physics determining the outcome of the merger, in addition to $M_{\rm TOV}$. For example, the mass ratio $q$ is known to play an important role in setting the angular momentum of the remnant [73]. Moreover, there might be peculiar features in the EOSs that influence the outcome of mergers, beyond $M_{\rm TOV}$. A more accurate determination of the separation between long- and short-lived remnants is beyond the scope of this work. For example, it would be interesting to reformulate the criterion for the formation of short- vs long-lived remnants probabilistically and to explore its crucial physical dependencies. In combination with estimates for $M_{\rm TOV}$, this would have the potential to provide novel constraints on the EOS of dense matter in the most extreme conditions reached in mergers. Unfortunately, the binaries included in the CoRe database were not selected randomly, and, in particular, some binaries are overrepresented (e.g., GW170817-like configurations), so our current numerical relativity data is not yet sufficient to realize this goal.
@@
We find that a relatively large value for the transition between long and short-lived NS remnants, $M_{\rm ls}\gtrsim 1.3M_{\rm TOV}$, is strongly favoured by the current data for a range of viable EoSs. This implies that the fraction of mergers resulting in the formation of long-lived remnants might extend over a broader mass range than traditionally expected. Motivated by these phenomenological results, we then searched for an indication of the long-short-lived transition in a subset of numerical simulations robust enough to resolve such regime change. While the transition region is found to be broad in the numerical data, there is clear evidence for a transition at $\sim 1.3M_{\rm TOV}$, in agreement with our findings.
Our results disfavour scenarios, such as certain types of QCD phase transitions, pion or kaon condensation, or other exotic physics that would lead to catastrophic pressure loss at several times nuclear density and temperatures of tens of MeV leading to the rapid gravitational collapse of binaries with total mass $M\lesssim 1.3M_{\rm TOV}$ [61]. That said, the formation of short-lived remnants with $M\gtrsim 1.3M_{\rm TOV}$ can still be accommodated if $M_{\rm TOV}\gtrsim 2.6M_{\odot}$, which however is in tension with constraints from GW170817 [4, 9, 10]. Our findings are consistent with the interpretation of the blue kilonova in GW170817 as a spiral-wave driven wind [74, 75, 76], or as the magnetized wind from a long-lived, but ultimately unstable magnetar [77, 78].
@@ Acknowledgements.
We thank Alejandra Gonzalez, Patrick Bush III, and Rossella Gamba for help with the CoRe database data; Floor Broekgaarden, Will Farr, Michela Mapelli, Soumendra Roy and Lieke van Son for informative discussions on the neutron star mass distribution, and Brian Metzger for very useful comments on our manuscript. We are especially grateful to Jillian Rastinejad and Andrew Levan for sharing their pre-publication data on the relative rates of long and short cbGRBs. RP acknowledges support from the National Science Foundation under Grant AST-2006839. OG is supported by the Flatiron Research Fellowship. The Flatiron Institute is supported by the Simons Foundation. ES was partially supported by NASA through Award Number 80NSSC21K1720. DR acknowledges support from the Sloan Foundation, from the U.S. Department of Energy, Office of Science, Division of Nuclear Physics under Award Number(s) DE-SC0021177 and DE-SC0024388, and from the National Science Foundation under Grants No. PHY-2020275, AST-2108467, PHY-2116686, and PHY-2407681.

### v2

@@ III.2 Long and Short GRBs from BNS mergers
Our analysis assumes that the VLNS population does not contribute to a sizeable fraction of cbGRBs for the reasons discussed earlier. However, as noted by Margalit et al. [66], some initially long-lived magnetars may evolve similarly to LLNSs, potentially powering an sbGRB. Therefore, we assess the implications of this scenario on our results by performing two additional simulations that consider the two extreme cases of the VLNS population contributing fully to either the short or the long cbGRB components, depending on whether the VLNSs collapse on a timescale of $\lesssim 1\,{\rm s}$ or $\sim 10\,{\rm s}$, respectively. We find that if VLNSs contribute exclusively to sbGRBs, the transition with $a=1.2$ is still inconsistent with the data for $M_{\rm TOV}<2.5M_{\odot}$, while the $a=1.25$ threshold is only allowed for $M_{\rm TOV}\gtrsim 2.4M_{\odot}$. The higher value $a=1.3$ has the best consistency in the 2.2-2.4 $M_{\odot}$ TOV range of masses, while higher values of $a$ would require softer EoSs closer to 2.2 $M_{\odot}$. Conversely, if VLNS contribute to the long cbGRB population only, we find that, within the 2.2-2.4 $M_{\odot}$ range of TOV masses, $a\lesssim 1.3$ is borderline consistent with the data, while both $a=1.35$ and $a=1.4$ have a broad range of consistency. These cases discussed above would constitute the most extreme departures from our fiducial model since either a smaller contribution from the VLNS population or a mixed contribution to both the long and short cbGRBs would make the final estimates closer to those shown in Fig. 4. Therefore, our results of a relatively high transition mass $M_{\rm ls}\gtrsim 1.3M_{\rm TOV}$ are quite robust within the limits and assumptions of the unified GRB model and the available observational data to date. A broad quantitative assessment of the robustness of our results with respect to the various model inputs is provided in the Appendix.
@@ III.3 Comparison with numerical relativity
Our results are summarized in Fig. 5. The data shows clear evidence for a transition around ${\sim}1.3{\color[rgb]{0,0,0}M_{\rm TOV}}$, consistent with the population synthesis results. The transition region is broad, spanning $1.3\lesssim{\color[rgb]{0,0,0}M_{\rm ls}}/M_{\rm TOV}\lesssim 1.4$, suggesting that there is additional physics determining the outcome of the merger, in addition to $M_{\rm TOV}$. For example, the mass ratio $q$ is known to play an important role in setting the angular momentum of the remnant [77]. Moreover, there might be peculiar features in the EOSs that influence the outcome of mergers, beyond $M_{\rm TOV}$. A more accurate determination of the separation between long- and short-lived remnants is beyond the scope of this work. For example, it would be interesting to reformulate the criterion for the formation of short- vs long-lived remnants probabilistically and to explore its crucial physical dependencies. In combination with estimates for $M_{\rm TOV}$, this would have the potential to provide novel constraints on the EOS of dense matter in the most extreme conditions reached in mergers. Unfortunately, the binaries included in the CoRe database were not selected randomly, and, in particular, some binaries are overrepresented (e.g., GW170817-like configurations), so our current numerical relativity data is not yet sufficient to realize this goal.
@@
We find that a relatively large value for the transition between long and short-lived NS remnants, $M_{\rm ls}\gtrsim 1.3M_{\rm TOV}$, is strongly favoured by the current data for a range of viable EoSs. This implies that the fraction of mergers resulting in the formation of long-lived remnants might extend over a broader mass range than traditionally expected. We remind the reader that our modeling relies on various input parameters, some of which are better constrained than others. While we adopted a fiducial model based on the best fits to current observational and numerical data from the literature, we have explored in the appendix how our conclusions would change if these input parameters deviated from their assumed values. Of particular importance is the dependence on the NS mass function for both the primary and secondary, which we expect to be more robustly calibrated in the future as GW and radio data of binary NSs grow.
Motivated by our phenomenological results, we then searched for an indication of the long-short-lived transition in a subset of numerical simulations robust enough to resolve such regime change. While the transition region is found to be broad in the numerical data, there is clear evidence for a transition at $\sim 1.3M_{\rm TOV}$, in agreement with our findings.
Our results disfavour scenarios, such as certain types of QCD phase transitions, pion or kaon condensation, or other exotic physics that would lead to catastrophic pressure loss at several times nuclear density and temperatures of tens of MeV leading to the rapid gravitational collapse of binaries with total mass ${\color[rgb]{0,0,1}M_{\rm tot}}\lesssim 1.3M_{\rm TOV}$ [66]. That said, the formation of short-lived remnants with ${\color[rgb]{0,0,1}M_{\rm tot}}\gtrsim 1.3M_{\rm TOV}$ can still be accommodated if $M_{\rm TOV}\gtrsim 2.6M_{\odot}$, which however is in tension with constraints from GW170817 [4, 10] and with multi-messenger NS data inference by [78]. Our findings are consistent with the interpretation of the blue kilonova in GW170817 as a spiral-wave driven wind [79, 80, 81], or as the magnetized wind from a long-lived, but ultimately unstable magnetar [82, 83].
@@ Acknowledgements.
We thank Alejandra Gonzalez, Patrick Bush III, and Rossella Gamba for help with the CoRe database data; Floor Broekgaarden, Will Farr, Michela Mapelli, Soumendra Roy and Lieke van Son for informative discussions on the neutron star mass distribution, and Brian Metzger for very useful comments on our manuscript. We are especially grateful to Jillian Rastinejad and Andrew Levan for sharing their pre-publication data on the relative rates of long and short cbGRBs. Last we thank an anonymous referee for insightful and constructive comments which helped to strengthen our work. RP acknowledges support from the National Science Foundation under Grant AST-2006839. OG is supported by the Flatiron Research Fellowship. The Flatiron Institute is supported by the Simons Foundation. ES was partially supported by NASA through Award Number 80NSSC21K1720. DR acknowledges support from the Sloan Foundation, from the U.S. Department of Energy, Office of Science, Division of Nuclear Physics under Award Number(s) DE-SC0021177 and DE-SC0024388, and from the National Science Foundation under Grants No. PHY-2020275, AST-2108467, PHY-2116686, and PHY-2407681.
Appendix: Dependence of results on the relevant input parameters
Here, we quantitatively examine how our predicted lbGRBs/sbGRBs ratio depends on the main input parameters. We vary one parameter at a time, exploring deviations from the fiducial model in each case. Since an increase or decrease in the lbGRBs/sbGRBs ratio causes all the curves in Fig. 4 to shift accordingly, we focus our analysis on a single reference value of $M_{\rm ls}/M_{\rm TOV}$, specifically $1.3$.
(a) Dependence on $M_{\rm sp}$
Fig. 6 compares the ratio lbGRBs/sbGRBs between our fiducial, low-limit case of $M_{\rm sp}/M_{\rm TOV}=1.15$, with the higher value $M_{\rm sp}/M_{\rm TOV}=1.2$ suggested in the literature [44]. The figure clearly shows that our fiducial choice of $M_{\rm sp}/M_{\rm TOV}=1.15$ is conservative with respect to our results: a higher value of the maximum mass of an uniformly rotating NS would strengthen the case of a high transition mass $M_{\rm ls}$.
Figure 6: Dependence of the ratio between sbGRBs and lbGRBs on the transition mass from VLNSs to LLNSs. The choice of $M_{\rm sp}/M_{\rm TOV}=1.15$ adopted here makes our results more robust with respect to the 1.2 value commonly adopted in the literature.
(b) Dependence on $M_{\rm th}$
While here we used the fit to $M_{\rm th}$ by Kashyap et al. [45] and Perego et al. [46], however several other fits exist in the literature (e.g. [84, 85, 86]), generally agreeing within a few percent of a solar mass. To evaluate the impact of variations in $M_{\rm th}$ on our results, we performed two additional sets of simulations: one where $M_{\rm th}$ was increased by a conservative $5\%$ of $M_{\odot}$ and another where it was decreased by the same amount. The small dependence on $M_{\rm th}$ derives from the presence of this variable in the expression for the disk mass fit. Figure 7 presents the results of this analysis, demonstrating that the variations induced by potential uncertainties in $M_{\rm th}$ are minor compared to the trends observed with varying $M_{\rm ls}$ in Figure 4.
Figure 7: Dependence of the ratio between sbGRBs and lbGRBs on the threshold mass for prompt collapse to a BH. A 5% $M_{\odot}$ variation around our adopted best fit to $M_{\rm thr}$ is considered, since various prescriptions from the literature generally agree within this range.
(c) Dependence on $M_{\rm d}$
As discussed in IIC, the disk mass adopted here is based on a fit produced from the combination of the results of a large number of numerical simulations. However, these are subject to various uncertainties, both of numerical nature as well as due to incomplete microphysics. Since these uncertainties are hard to exactly quantify, here we adopt an inverse approach: we study by what amount the fit to the disk mass needs to be incorrect for our results to weaken.
First we note that, if the disk mass was systematically underestimated, then our results would be even more robust towards high values of $M_{\rm ls}$, since the lbGRBs/sbGRBs ratio would shift upwards. Therefore, we present numerical results only for the scenario in which the disk mass is systematically overestimated by the fit. Figure 8 illustrates how the ratio for the $M_{\rm ls}=1.3$ case would change if the disk mass were $90\%$, $70\%$, or $50\%$ of the fitted value. As shown in the figure, for our conclusions to be significantly weakened, the disk mass would need to have been overestimated in the current fits by a factor of about 2.
Figure 8: Dependence of the ratio between sbGRBs and lbGRBs on the mass of the disk around the post-merger object, making the assumption that the disk mass is systematically lower than what provided by the best fit, by various amounts. Higher values of the disk mass would increase the ratio and make our conclusions more robust.
(d) Dependence on the NS mass distribution
Last we perform further simulations to explore possible deviations from the assumed NS mass function. This exploration is motivated by the fact that observational NS samples are heterogeneous, and could be biased by selection effects (see e.g. [87]). Ideally, the NS mass function (NSMF) as input for this analysis would be one calibrated on the subset of NS-NS systems which merge in the Hubble time (sample gathered from a combination from GW observations and radio surveys). However, until this subset is large enough, we can only work with the full sample, with the understanding that it may or may not be a faithful representation of the merging one.
Figure 9: Dependence of the ratio between sbGRBs and lbGRBs on variations in the NS mass distribution wit respect to our adopted best fit. The top panel explores the dependence on the relative strengths $A_{1}$ and $A_{2}$ of the low- and high-mass peaks of the NS mass distribution, respectively. In the lower panel the NS mass distribution of the primary is taken from the best fit, but, for exploratory purposes, the distribution of the mass ratio $q$ is assumed either flat between 0.5 and 1, or a relatively narrow one-sided Gaussian with maximum at 1 and dispersion $0.15$.
For this analysis, we continue to adopt a bimodal NSMF, since this has been a robust finding with the latest NS samples by several independent researchers (e.g., [35, 36, 37, 38, 39, 40]), and it has been suggested to hold independently also for the GW subsample [39]. However, we explore the impact of varying the relative fraction of low-mass and high-mass NSs on our results. To this end, we conducted several simulations, gradually adjusting the normalizations $A_{1}$ and $A_{2}$ of the two peaks—from their fiducial values of 0.54 and 0.46, respectively—until noticeable deviations from the fiducial model emerged.
The results are presented in the top panel of Fig. 9, where the normalizations of the two NS populations are each increased or decreased by a factor of 0.2, corresponding to a variation of approximately $\sim 40\%$ relative to the fiducial values. If the high-mass population were more dominant, our conclusions would be reinforced, as the fractions of SLNSs and pcBHs would increase. Conversely, a relatively larger fraction of lower-mass NSs would reduce the occurrence of these remnant types. However, as shown in the figure, a substantial deviation from the fiducial (best-fit) values is required before our conclusions begin to weaken.
Lastly, we examine the impact of the mass ratio $q$. In our fiducial model, we adopted a random pairing assumption, generating both the primary and secondary masses from the same initial mass function. This resulted in a distribution peaked around $q=1$ but with a broad tail. To explore how variations in $q$ affect the results, we consider two extreme models: one in which $q$ is drawn from a flat distribution, and another in which it follows a Gaussian distribution centered at $q=1$ with a standard deviation of $\sigma=0.15$, which significantly favors systems with comparable masses relative to the fiducial model.
The effect of $q$ in our simulations is twofold. For a given primary mass distribution, lower values of $q$ correspond to lower secondary masses, leading to a decrease in $M_{\rm tot}$. This, in turn, reduces the fraction of SLNS and pcBH remnants, thereby lowering the lbGRB/sbGRB ratio. Conversely, lower values of $q$ tend to produce higher disk masses, which counteracts the impact on the lbGRB/sbGRB ratio. Both of these effects are highly sensitive to the EoS, leading to non-monotonic trends in their influence on the results. For the flat $q$ distribution, these opposing effects roughly balance each other, except at low values of $M_{\rm TOV}$, where the system is more sensitive to reductions in $M_{\rm tot}$. On the other hand, a $q$ distribution that strongly favors nearly equal-mass systems would be primarily influenced by the corresponding decrease in disk mass, resulting in a lower yield of lbGRBs compared to sbGRBs. Ultimately, future observations of binary neutron stars through gravitational waves and radio surveys will be crucial in refining the neutron star mass function for both the primary and secondary components (if they differ). This will provide a more robust foundation for our analysis.
(e) Dependence on the observational data
If the accumulation of future data were to alter the current lbGRB/sbGRB ratio, it would undoubtedly impact our inferences on $M_{\rm ls}$. The direction of this effect is evident in Fig. 4 (as well as in all the figures in the appendix). An increase in the relative number of long cbGRBs would weaken our conclusions, whereas a rise in the relative number of short cbGRBs would strengthen them.